OpenStudy (anonymous):

Find four consecutive integers, such that the product of the first and second integers minus 2 is equal to 3 times the product of the third and fourth integers.

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

help please

OpenStudy (mertsj):

Do you know how to represent 4 consecutive integers?

OpenStudy (anonymous):

-6 -5 -4 -2?

OpenStudy (anonymous):

-3**

OpenStudy (mertsj):

No. x x+1 x+2 x+3

OpenStudy (mertsj):

Now using that, you can write the equation.

OpenStudy (anonymous):

(x+1)(x+2)(x+3)

OpenStudy (anonymous):

x(x+1)(x+2)(x+3)

OpenStudy (mertsj):

What is the product of the first and second ones?

OpenStudy (anonymous):

-1, -2 ?

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

x=first of four consecutive integers x+1= second of four consecutive integers x+2= third of four consecutive integers x+3 = fourth of four consecutive integers

OpenStudy (mertsj):

What is the product of the first and second of the four consecutive integers?

OpenStudy (anonymous):

I dont get it :( @Mertsj

OpenStudy (anonymous):

can you show it to me please :( ?

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

@partygirl do you know how to solve this?

OpenStudy (anonymous):

-2, -1,0,1 or -5,-4,-3,-2

OpenStudy (anonymous):

how would I explain though @Arddhendu

OpenStudy (anonymous):

Let the consecutive integers be x, x+1, x+2, x+3 As per the question x(x+1)-2=3(x+2)(x+3) x^2+x-2= 3x^2+15x+18 2x^2+14x+20=0 2(x^2+7x+10)=0 x^2+7x+10=0 x^2+2x+5x+10=0 x(x+2) +5(x+2)=0 (x+2)(x+5)=0 so, x+2=0 or x+5=0 so, x= -2 or -5 If x =-2, then the 4 consecutive integers are -2, -2+1, -2+2, -2+3 i.e., -2, -1, 0,1 If x=-5, then the 4 consecutive integers are -5, -5+1, -5+2, -5+3 i.e., -5, -4, -3,-2

OpenStudy (anonymous):

is that everything ? @Arddhendu

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok Im going to write it exactly how you wrote it

OpenStudy (anonymous):

thank you so much ! @Arddhendu